OpenStudy (anonymous):

What is x when 4(3x+5)

jigglypuff314 (jigglypuff314):

was the value "4(3x + 5)" set equal to something?

OpenStudy (anonymous):

No I just need to find out what x is

jigglypuff314 (jigglypuff314):

what was the topic involved? graphing or solving or other?

OpenStudy (anonymous):

Just solving i think it was on a mock exam

OpenStudy (skullpatrol):

4(3x + 5) is a variable expression. You can't solve for anything.

OpenStudy (anonymous):

Well yeah that's where I got stuck because i can't work it out so what would I have to put on the paper

jigglypuff314 (jigglypuff314):

hmm, from what you are implying might I ask what the original question was? or was it just " 4(3x + 5)"

OpenStudy (skullpatrol):

Is this the exact question? What is x when 4(3x+5)

OpenStudy (anonymous):

Yet that was the question what is x in 4(3x+5)

OpenStudy (anonymous):

Yeah that's the question

OpenStudy (skullpatrol):

You could say x is the variable that is multiplied by 12 after distributing the 4 : 12x + 20

OpenStudy (anonymous):

Sorry forgot it =28

OpenStudy (skullpatrol):

>8(

OpenStudy (anonymous):

If that's what you meant

OpenStudy (skullpatrol):

That is EXACTLY what we mean!!!

jigglypuff314 (jigglypuff314):

see, that was what I was asking before xD so 12x + 20 = 28 subtract 20 from both sides divide both sides by 12

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

I don't get ya

jigglypuff314 (jigglypuff314):

12x + 20 = 28 subtract 20 from both sides -20 -20 12x = 8 divide both sides by 12 /12 /12 x = 8/12 then simplify

OpenStudy (skullpatrol):

$$\Huge 4(3x+5)=28$$

OpenStudy (skullpatrol):

Use the idea of balance to solve the equation.

OpenStudy (skullpatrol):

Because errors may occur in solving equations, you should check that your solution of the final equation satisfies the original equation.

OpenStudy (anonymous):

Ok thank you very much this has really helped.

OpenStudy (skullpatrol):

Thanks for asking and trying to understand :)